Session: Feb 2001 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Orfei at Sat Feb 24 09:15:32 MET 2001) Success

Jodrell_Bank (Alastair Gunn at Sun Feb 25 12:59:22 MET 2001) Minor failures
Problems with vacuum failure on second tape (at 00:54). Some loss of data until 02:40 when problem seemed to be resolved. Extreme RFI especially in VCs 3 and 4. No other known problems.

Onsala (A. Polatidis at Mon Feb 26 02:07:33 MET 2001) Minor failures
Lost vacuum on second tape at 5:45UT. The tape got mangled and the operator did not manage to restart the experiment. Lost the last 2hrs 15min.

Torun (K.Borkowski at Mon Feb 26 10:40:46 MET 2001) Success
Essentially a success, however the first two tape passes (forward and reverse) on the first tape were recorded with different tension (controlled on the rear of the front panel of recorder). The original tension seemed to be inappropriate for this tape and speed, causing a bad packing (however not that very bad as we have experienced in the past).

Westerbork (Tony Foley at Tue Feb 27 11:52:03 MET 2001) Minor failures
Done, as requested, in single dish mode with RT7. Large sections were outside hour angle range, so could not be observed 14:00-17:16UT and 05:13-05:16UT. Some EMI.

Effelsberg (A. Kraus at Tue Feb 27 12:21:25 MET 2001) Minor failures
Large parts have been lost because of snowfall: UT1427-1506, UT2021-2300, UT0130-0400, UT0430-0530, UT0730-end.

Noto (C.Trigilio at Wed Feb 28 15:40:19 MET 2001) Success